OpenStudy (anonymous):

Use the substitution method to solve the following system of equations x + 2y – z = 2 3x + y – 2z = –15 3x + 4y – 3z = –2

OpenStudy (anonymous):

solve the first equation for x. x = 2-2y+z then substitute x into the next equation and solve for y 3*(2-2y+z) +y - 3z = -15 then substitute that in to the third equation and solve for z

OpenStudy (anonymous):

oops.. typo. x = 2-2y+z then substitute x into the next equation and solve for y 3*(2-2y+z) +y - 2z = -15 then substitute that (after you solve for y) in to the third equation (along with what you have for x) and solve for z

OpenStudy (anonymous):

so what solution do i get